The left and right page numbers of an open book are two consecutive integers whose sum is 337 find These page numbers smaller & larger page.
Answer by ikleyn(52946)

.

Let n be the left page number; then the right page number is (n+1).


You are given


    n + (n+1) = 337.


Hence,


    2n = 337-1 = 336,

     n = 336/2 = 168.


ANSWER.  The pages are  168 and 169. 

Solved.